• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M. Cummings

    Log entry time 08:00:22 on March 29, 2012

    Entry number 368013

    keyword=Owl Shift Summary

    Owl Shift Summary
    SL: Melissa Cummings
    TO: Mikhail Yurov 
    3rd: Jechiel Lichtenstadt 
    00:00 Start of shift, hall is in controlled access to remove sieve slit.
    00:10 Sieve is OUT, going back to beam permit.
    00:15 Alaexandre starts his BCM calorimeter runs
    00:35 Start target polarization
    01:25 Target polarized to 40%, calling for 50nA
    01:30 RHRS has very low rate (~2kHz) compared to ~100kHz on the LHRS.  dp
    variables look similar on left/right.  Beam position is still centered.    
    03:30 Called for escorted access to check that sieve slit is actually
    out.  It seems the Right sieve slit was in fact still in.
    03:50 Back in Beam Permit, asked for 50nA.  RHRS rate seems back to
    normal, start taking production data.
    05:30 Done with elastic setting. Taking dilution runs.
    06:30 Done with dilution runs.  Change L&R HRS dipole settings to 2.072.
    Change Left Septum current to 661.3 A and Right Septum Current to 735.1 A
    07:15 MCC is tuning beam into Hall B, will be down for a little while.
    08:00 End of Shift, still waiting for MCC to finish tuning into Hall B.
    Ready to start taking data for p=2.072
    Runs Taken:
    p = 2.228
    3440/40488              HWP IN
    3444/22477,22478/40490  HWP OUT, Beam trip at end of run
    3445/22479/40491        HWP OUT
    3446/22480/40492        HWP OUT
    Dilution Runs:
    Carbon w/ He, Slow Raster on:  3447/22481/40493
    Dummy w/ He,  Slow Raster on:  3448/22482/40494
    Empty w/ He,  Slow Raster on:  3449/22483/40495
    Carbon w/ He, Slow Raster off: 3453/22487/40500       
    Dummy w/ He,  Slow Raster off: 3452/22486/40499      
    Empty w/ He,  Slow Raster off: 3451/22485/40498